OpenStudy (anonymous):

y=2x^2+4x-8 The axis of symmetry is the line x = ____.

OpenStudy (anonymous):



OpenStudy (jdoe0001):

find the vertex for it, usually you could find it with \(\bf \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\) so the "x" coordinate off the pair, will be the value for the axis of symmetry

OpenStudy (anonymous):

what is the answer i will give you a meadl @jdoe0001

OpenStudy (jdoe0001):

well, what is it? yeah, well, you tell me :) we help by clearing out some things folks may be off about not by getting answers so they can transcribe them

OpenStudy (jdoe0001):

I mean, using the \(\bf \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\) can't get simpler than that

OpenStudy (cwrw238):

you can also find the answer by converting the equation in vertex form

OpenStudy (cwrw238):

y=2x^2+4x-8 y = 2(x^2 + 2x - 4) = 2(x + 1)^2 - 5) = 2 (x + 1)^2 - 10 this is the vertex form te value of x corresponding to the vertex is when x + 1 = 0 - now you can find the line of symmetry

OpenStudy (jdoe0001):

hmmm, I'd not bother, he just wanted it done for him

OpenStudy (cwrw238):

yea i know but i'm hoping i might cajole him into trying to do something for himself

OpenStudy (cwrw238):

he's gone anyway!!!!

OpenStudy (jdoe0001):

if he's not going to use a simple equation like the one I posted, very unlikely he'd do a square completion, some folks just misundertand the meaning of "help", sadly

OpenStudy (cwrw238):

yea - i guess you are right - i just thought i'd try...